OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (aripotta):

x = -1, x = -5

OpenStudy (anonymous):

\[−3|2x + 6| = −12\]divide both sides by -3 to isolate the absolute value:\[|2x+6|=4\]now set up two equations that account for both negative and positive, as so:\[2x+6=4~~~~~~~~~~2x+6=-4\] and now solve for both :)

OpenStudy (aripotta):

o-o this was a month closed bruh

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i didnt even see that